OpenStudy (anonymous):

What is the second term of (5v + s)^5 ?

OpenStudy (anonymous):

Are you using Pascal's Triangle? Or are you using combination theory?

OpenStudy (anonymous):

5 C 1 will be 5 , so you will have an automatic factor of 5. Then you have (5v)^4 times s. So, you have (5)[(5v)^4](s) = (5)(5^4)(v^4)(s) = (5^5)(v^4)(s)

OpenStudy (anonymous):

Making sense to you or do you want to go over any part of it?